OpenStudy (anonymous):

PLEASE HELP IF YOU CAN! √3 x √21

OpenStudy (anonymous):

oops \[\sqrt{3\times 21}=\sqrt{63}=\sqrt{9\times 7}=\sqrt{9}\times \sqrt{7}=3\sqrt{7}\]

OpenStudy (anonymous):

\[3\sqrt{7}\]

OpenStudy (anonymous):

How did you get 9x7?

OpenStudy (anonymous):

9 and 7 and just factors of 63...he used them because 9 is a perfect square that he can take the root of

OpenStudy (anonymous):

that's what i thought, thank you. Now do you have to find a perfect square in order to use it as an equation that results in the number 63?

OpenStudy (anonymous):

mmm well \[\sqrt{63 }\] and 3\[\sqrt{7}\] are the same answer... one is just simplified. but yes when multiplying rads just multiply the two numbers under the radical and see if there are any perfect suqares that are factors of that number

OpenStudy (anonymous):

Thank you kind sir :)